OpenStudy (anonymous):

help!?!? What is the weight (in grams) of a liquid that exactly fills a 163.5 milliliter container if the density of the liquid is 0.492 grams over milliliter? Round to the nearest hundredth when necessary and only enter numerical values, which can include a decimal point.

OpenStudy (anonymous):

\[Density = \frac{ Mass }{ Volume }\] \[0.492\ \frac{ g }{ mL } = \frac{ x\ grams }{ 163.5\ mL }\]

OpenStudy (anonymous):

can you solve for x?

OpenStudy (anonymous):

i dont know how

OpenStudy (anonymous):

when you have: \[\frac{ A }{ B } = C\] to solve for A, you must multiply both sides by B. \[\frac{ A }{ B } \times B = C \times B\]\[A = BC\] similarly, to isolate x and get its value, we must multiply both sides by 163.5 \[x = 0.492 \times 163.5\]

OpenStudy (anonymous):

x=80.442

OpenStudy (anonymous):

yes. and they say "round to the nearest hundredth" the hundredth place is the second place after the decimal point so we only keep 2 of the 3 digits

OpenStudy (anonymous):

so the answer will be 80.44

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

glad to help ^_^